# FedOC: Optimizing Global Prototypes with Orthogonality Constraints for Enhancing Embeddings Separation in Heterogeneous Federated Learning

### v1

@@ 5.5 Robustness to Statistical Heterogeneity
To evaluate the robustness of all baseline methods under different non-IID data distributions, we conducted experiments on two dataset configurations. As demonstrated in Tab. 4, FedOC ensures optimal inter-class separation by leveraging orthogonality constraints, reducing embeddings overlap even in highly skewed data distributions. This demonstrates its effectiveness in promoting generalization in statistical heterogeneity environments.
@@ 5.6 Impact of Feature Dimension
To evaluate the robustness of different algorithms across feature dimensions $K$, we conducted experiments with three feature dimensions: 128, 256, and 1024, as dipicted in Fig. 4. Results indicate that all baselines maintain a high degree of robustness, showing minimal variation in performance.
@@
Figure 4: Test accuracy with different feature dimension
@@ 5.7 Ablation Study
This section analyzes the ablation study in Tab. 5. FedOC without orthogonality constraints (FedOC w/o OC) aligns the embedding layer with the global prototype derived from weighted averaging. While the performance on CIFAR-10 is passable, FedOC w/o OC exhibits a sharp performance decline on CIFAR-100 and Flowers102 datasets, likely due to substantial global prototype overlap. Introducing the orthogonality constraints in FedOC markedly enhances prototype separation, improving performance by explicitly increasing angular separation across classes. FedOC outperforms FedTGP, as reducing prototype similarity integrates more effectively with cross-entropy (CE) loss than merely increasing prototype distances.
Table 5: Test the accuracy (%) on Cifar100 of FedProto, FedOC w/o orthogonality constraints and FedOC.
FedTGP FedOC w/o OC FedOC
Cifar10 81.60 ± 0.30 78.05 ± 0.52 85.71 ± 0.10
Cifar100 52.17 ± 0.74 35.11 ± 0.63 61.12 ± 0.15
Flowers102 52.43 ± 0.64 42.43 ± 0.57 64.15 ± 1.66
@@ 5.8 Hyperparameter Combination
Table 6: Test the accuracy (%) on Cifar100 of FedOC with various parameters combination.
@@ 6 Conclusion
In this work, we propose a novel HtFL method, FedOC, which employs orthogonality constraints on global prototypes at the server side to improve inter-prototype separation and maximize directional independence while preserving semantic integrity. On the client side, FedOC aligns the embeddings with prototypes, enhancing its compatibility with CE loss. This approach overcomes the limitations of FedTGP by enhancing prototype separation through angular divergence rather than merely increasing Euclidean distances between prototypes. Extensive experiments indicate that FedOC surpasses baseline methods, achieving superior accuracy and robustness across a range of heterogeneous scenarios.

### v2

@@ 5.5 Robustness to Statistical Heterogeneity
To evaluate the robustness of all baseline methods under different non-IID data distributions, we conducted experiments on two dataset configurations. As demonstrated in Tab. 4, FedORGP ensures optimal inter-class separation by leveraging orthogonality regularization, reducing embeddings overlap even in highly skewed data distributions. This demonstrates its effectiveness in promoting generalization in statistical heterogeneity environments.
@@ 5.6 Impact of Feature Dimension
To assess the robustness of all baselines to the $K$, we performe experiments with three feature dimensions: 128, 256, and 1024, as indicated in Fig. 4. Results indicate that all baselines maintain a high degree of robustness, showing minimal variation in performance.
@@
Figure 4: Test on Cifar-100 (10 classes per client) under different feature dimensions with HMG8.
@@ 5.7 Ablation Study
This section analyzes the ablation study in Tab. 5. FedORGP without orthogonality regularization (FedORGP w/o OC) aligns the embedding layer with the global prototype derived from weighted averaging. While the performance on CIFAR-10 is passable, FedORGP w/o OC exhibits a sharp performance decline on CIFAR-100 and Flowers102 datasets, likely due to substantial global prototype overlap. Introducing the orthogonality regularization in FedORGP markedly enhances prototype separation, improving performance by explicitly increasing angular separation across classes. FedORGP outperforms FedTGP, as reducing prototype similarity integrates more effectively with cross-entropy (CE) loss than merely increasing prototype distances.
Table 5: Test on three datasets under the practical setting using the HMG8 for ablation study.
FedTGP FedORGP w/o OR FedORGP
Cifar-10 84.75 ± 0.06 78.05 ± 0.52 85.71 ± 0.10
Cifar-100 53.12 ± 0.24 35.11 ± 0.63 61.12 ± 0.15
Flowers102 58.60 ± 0.14 42.43 ± 0.57 64.15 ± 1.66
@@ 5.8 Hyperparameter Combination
Table 6: Hyperparameter search on Cifar-100 with HMG8.
@@ 6 Conclusion
In this work, we propose a novel HtFL method, FedORGP, which employs orthogonality regularization on global prototypes at the server side to improve inter-prototype separation and maximize directional independence while preserving semantic integrity. On the client side, FedORGP aligns the embeddings with prototypes, enhancing its compatibility with CE loss. This approach overcomes the limitations of FedTGP by enhancing prototype separation through angular separation rather than merely increasing Euclidean distances between prototypes. Extensive experiments demonstrate that FedORGP consistently outperforms all baseline methods, and exhibits superior performance in scenarios where model and statistical heterogeneity coexist.